OpenStudy (anonymous):

Help please. (: Medal. A point on the rim of a wheel moves with a velocity of 100 feet per second. Find the angular velocity of the point if the diameter of the wheel is 8 feet. 12.5 rad/sec 25 rad/sec 400 rad/sec 800 rad/sec

OpenStudy (anonymous):

dia = 8 ft radius= 4 ft angular velocity = V/r = 100/4 = 25 rad/sec b. 25 rad/sec

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

Yes thank you so much!

OpenStudy (anonymous):

no problem